Proble 9 Part A What is molarity of a solutic that contains 500g of vitamin B1 hydrochloride (molar mass 337 g/mol in 180 mL of solution? a molant My Answers Gve up
Answer #1

A)

olarity = mol / volume

mol = mass/MW

mol= 50 / 337 = 0.148367 mol

V = 0.180 L

M = 0.148367/0.180

M = 0.82426 M